A drug tester claims that a drug cures a rare skin disease

75​%

of the time. The claim is checked by testing the drug on

100

patients. If at least

68

patients are​ cured, the claim will be accepted.

Find the probability that the claim will be rejected assuming that the​ manufacturer's claim is true. Use the normal distribution to approximate the binomial distribution if possible.

The probability is

nothing.

​(Round to four decimal places as​ needed.)
Answer #1

Sample size, n = 100

P(a patient getting cured), p = 0.75

q = 1 - p = 0.25

Mean = np

= 100x0.75

= 75

Standard deviation = \sqrt{npq}

= \sqrt{100\times 0.75\times 0.25}

= 4.33

P(X < A) = P(Z < (A - mean)/standard deviation)

P(test claim will be rejected) = P(less than 68 patients are cured)

= P(X < 67.5) (continuity correction is applied)

= P(Z < (67.5 - 75)/4.33)

= P(Z < -1.73)

= 0.0418